1. A single cylinder 1.6 lt, spark ignition engine is operating on the ideal Otto cycle at a compression ratio of 9. The ambi2. Design a turbocharger to boost the inlet air pressure of the engine given in Question 1. This device consists of an exhausSoin: - fuel consumption rate ( m) = 131% = 13 x16?m? volume of cylinder (V) = 1:64 = 16x16?m? compression ratio (CR) = 9. inf2 = P₂ (CR)* (In I) & max, Iman P, = 1x1944 → Po = 21.674 bar 2 2 Process 2-3, constant volume heat addition: Vg = V2=0-148b.) Net work output, Wret = HA-HR = 127.349-52.296 Whet 75.053 kJ Ans Engine output in hop- Ihp = 0.7457KIS 75.053 3. Wnet 0.

The solution belongs to the 1st question, i need the solution for the 2nd question.
Answer #1

The solution is shown below if any doubts please post in comment section.

02). Ty = 108 l. 413k., Pilbar Ti - 293k. 15 = 2bar A 5 5 S $ Torbo Charger turbine РЧ T4 11 V P. Pup = P1 1081.43 (1) - 3.7exit pressure ne = Ti-To T - TO .85 = T1 - 293 445.77-293 So Ti - 492.8545k = Pi РО To 004 [isen tropic temperature] 1.4 apolRy 8.734 bar = P4 1081.42 - T4 15 Ty N 14-1 4 80734 T5 1081.43 To= 709, 72K n=0.85 = T4-T51 T4-T5 T5l = Tu - 185 [14-15] 1081ulturbine - (oria tmp) C C Tu-15] (0.0896 + 2.2024103) (1.005) [1081.43 - 759.5265] Band iteration - 29.7 kW vane] 29.77 = Wc